# Efficient Data Generation for Source-grounded Information-seeking Dialogs: A Use Case for Meeting Transcripts

### v2

@@ 1 Introduction
In summary, our main contributions are: (1) we explore the feasibility of using LLMs for data generation in source-grounded information-seeking dialogs, demonstrating improved overall quality and efficiency compared to a fully manual approach; (2) we release the MISeD dataset [^1] We make our datasets publicly available. – the first dialog dataset over meeting transcripts, consisting of verified source-grounded dialogs with transcript attribution, as well as a corresponding smaller fully manual (WOZ) dataset; (3) we introduce baseline models, an evaluation procedure, and baseline results for meeting-grounded dialog tasks, through which we assess our proposed approach and the MISeD dataset.
@@ 4.2 Dialog Annotation and Validation
Following automatic dialog generation, trained annotators assessed the generated queries and responses while identifying supporting attribution spans within the source text. We employed 3 professional annotators (details in Appendix B) and conducted a pilot study measuring inter-annotator agreement over 525 dialog turns (B.0.1). The high level of agreement observed confirmed the annotation task’s clarity and criteria consistency, allowing us to proceed with a single annotator annotating each dialog in the main phase.
@@ 4.2.1 Query assessment
Annotators verify query validity by answering the question: "Is the question understandable and makes sense?". If a query is marked as invalid, it is removed from the dialog. We also remove all subsequent dialog turns to avoid cascading errors. Valid queries are then annotated with metadata, including their type (‘general’, ‘specific’, or ‘yes/no’) and whether they depend on prior turns (‘context-dependent’ or not).
@@ 4.2.2 Response annotation: attribution, validation, and editing
Due to the current limitations of LLM-based attribution detection (which we also observed in §7.3.2), we opted for fully manual attribution annotation. For each generated response, annotators identify supporting text spans in the transcript while also verifying that the response is factually correct and editing the response as necessary to ensure its accuracy (more details in Appendix B). If a query cannot be answered from the transcript, it is marked as ‘unanswerable’ and annotators ensure the response accurately conveys the lack of information.
@@ 5 Datasets
This section describes the MISeD dataset (§5.1), constructed in our semi-authomatic method (§4). We also present a smaller independent dataset created using the fully manual Wizard-of-Oz methodology (§5.2), created to assess whether MISeD data is comparable to human-generated data.
@@ 5.1.2 Dataset statistics
To assess the quality of the MISeD responses, we manually rated a sample of them, suggesting that the quality of MISeD responses is comparable to fully-manual responses (Appendix C).
@@ 5.2 Wizard-of-Oz Dataset
The collected WOZ data comprises 70 dialogs based on meetings from the test split, with a total of 700 query-response pairs. This data is used in subsequent evaluations (Section 7) for a comparison of model performance.
@@ 6.1 Response Quality Evaluation
Agent responses are evaluated against the gold responses in the test dataset.
@@ 6.1.1 Human evaluation
We conduct human evaluation on a random subset of 100 queries. The annotator is presented with the current user query, the gold response and the model responses, as well as the meeting transcript and the dialog history. To quantify the content overlap between the model response and the gold response, annotators provide scores for recall (how much of the gold response is covered by the model response), and precision (how much of the model response is covered by the gold response). Both scores are on a Likert scale of 1 to 4, corresponding to ‘Not Covered’, ‘Slightly Covered’, ‘Mostly Covered’ and ‘Fully Covered’.
@@ 6.2.2 Automatic evaluation
For automatic attribution evaluation we use the NLI model TrueTeacher (Gekhman et al., 2023) to approximate human entailment judgement, automatically computing recall (the proportion of response sentences fully entailed by the attributions) and precision (the proportion of attributions which contribute to the entailment of the response). Appendix D details the method implementation and discusses its limitations.
@@ 7 Baseline Models and Results
Human scores Automatic scores
@@
Human scores Automatic scores
@@ 7.1 Models
Semi-automatic (n=2922) 1.91 1.97 29.73 10.90 23.64 0.40 0.28 0.25 0.26
Semi-automatic (n=500) 1.93 1.99 30.52 11.10 24.10 0.40 0.26 0.16 0.20
Fully Manual (n=500) 1.91 1.94 31.88 12.64 26.17 0.41 0.02 0.36 0.04
Table 4: Evaluation using different training sets. Results on 200 WOZ test examples, using Gemini Pro finetuned.
@@ 7.3.1 Response quality
(3) The scores on WOZ test set exhibit the same trends, but are lower for all models. The lower performance for the non-finetuned models suggests that fully-manual WOZ data is more challenging, being created freely by human annotators (vs. the templated methodology of MISeD). As expected, the finetuned models performed better on the MISeD test set, as it was created by the same protocol as the training data.
@@ 7.3.2 Attribution quality
Observing the more reliable human scores, we see that the zero-shot models suffer from low recall, due to their tendency to provide attribution less frequently. Finetuning with MISeD data improves performance, substantially increasing recall with a small drop in precision, notably increasing F1 scores for both test sets. The automatic scores show similar though weaker trends, but we regard them as less reliable (see Appendix D).
@@ 7.4 Comparison of Training Sets
In this section we compare the performance of models trained on our semi-automatic MISeD data to those trained on the manual WOZ data. We split our WOZ dataset (700 dialog turns) into training and test sets (500:200, respectively). We compare training with (i) 500 WOZ examples (ii) 500 semi-automatic MISeD examples (iii) the full MISeD training set (2922 examples). For response generation, Table 4 suggests that training on 500 semi-automatic MISeD examples yields comparable results to training on the same amount of fully manual examples. For attribution, the model benefits from a larger number of training examples, suggesting that 500 examples are insufficient for this subtask. Notably, the model trained on WOZ only predicted attributions for 6 out of 200 test examples; thus we didn’t perform a manual attribution evaluation (as in Table 3). To summarize, training on our semi-automatic MISeD data yields response generation quality that is on par with training on the fully manual WOZ data, while also improving attribution, and saving time and effort.
@@ Appendix B Annotation details
B.0.1 Inter-annotator agreement
The held-out pilot data was annotated by all three annotators. Each dialog turn was evaluated (as a yes/no question) by each annotator based on each of the following criteria, for which we calculated Cohen’s Kappa agreement:
  1. 1.
    “Is the question understandable and makes sense?” (Kohen’s Kappa for each pair of annotators 1.0, 1.0, 1.0).
  2. 2.
    “Does the meeting transcript contain the right answer?” (Kohen’s Kappa for each pair of annotators: 1.0, 0.93, 0.93)
  3. 3.
    “Is the provided answer correct based on the meeting transcript?” (Kohen’s Kappa for each pair of annotators: 0.86, 0.90, 0.80).
Annotators were asked to edit the model responses in case they replied “yes” to the first two aforementioned questions (“Is the question understandable and makes sense?”, “Does the meeting transcript contain the right answer?”).
The high level of inter-annotator agreement suggests that the annotation task is well-defined and the criteria are consistently interpretable. Therefore we proceeded with a single annotator for the main annotation phase.
Appendix C MISeD Response Quality Assessment
MISeD dialogs were generated by an LLM and then validated and corrected by humans (§4). To assess the quality of the responses created through this semi-automatic process, we aimed to compare them with responses created fully-manually. To that end, we collected human responses to a randomly selected sample of 100 MISeD queries. Annotators were provided with the meeting transcript and the dialog history up to the selected query and were tasked with answering the query.
Next, we provided a new group of annotators with these manually-generated responses, alongside the original MISeD response for the each query (in randomized order, along with the corresponding query, dialog history, and transcript). Annotators assessed which response was overall superior, considering correctness, grounding, and clarity, using a scale of ‘Equally Good’, ‘Slightly Better’, or ‘Substantially Better’. Table 6 presents the results.
Overall, MISeD responses were ranked as better in 55 pairs, compared to 43 pairs for fully-manual responses. A statistical sign test yields $P(x\geq 55)=0.13$ ($H_{0}:p=0.5$; $n=98$). These results suggest that the quality of MISeD responses might be better than fully-manual responses, but a larger sample would be needed to confirm these findings and establish statistical significance.
Response evaluation categories Count Sum
MISeD ‘substantially better’ 29 55
MISeD ‘slightly better’ 26
‘equally good’ 2
Human-only ‘slightly better’ 21 43
Human-only ‘substantially better’ 22
Appendix D Automatic Evaluation of Attribution Implementation
@@
Model Rouge-1 Rouge-2 Rouge-L BLEURT
Gemini Pro 27.82 10.08 22.03 0.38
Gemini Pro Finetuned on MISeD 30.31 11.39 24.26 0.40
Gemini Pro Finetuned on QMSum 27.72 8.89 21.60 0.38
Table 13: Average scores for response evaluation on the WOZ test sets.

### v3

@@ 1 Introduction
In summary, our main contributions are: (1) we explore the feasibility of using LLMs for data generation in source-grounded information-seeking dialogs, demonstrating improved overall quality and efficiency compared to a fully manual approach; (2) we release the MISeD dataset [^1] We make our datasets publicly available at https://github.com/google-research-datasets/MISeD. – the first dialog dataset over meeting transcripts, consisting of verified source-grounded dialogs with transcript attribution, as well as a corresponding smaller fully manual (WOZ) dataset; (3) we introduce baseline models, an evaluation procedure, and baseline results for meeting-grounded dialog tasks, through which we assess our proposed approach and the MISeD dataset.
@@ 4.2 Dialog Annotation and Validation
Following automatic dialog generation, trained annotators assessed the generated queries and responses while identifying supporting attribution spans within the source text. We first employed 3 professional annotators (details in Appendix B) and conducted a pilot study, measuring Cohen’s Kappa agreement between each pair of annotators, over 525 dialog turns. Agreement scores are presented in Table 1.
Assessment Question Pair A Pair B Pair C
“Is the question understandable and makes sense?” 1.0 1.0 1.0
“Does the meeting transcript contain the right answer?” 1.0 0.93 0.93
“Is the provided answer correct based on the meeting transcript?” 0.86 0.90 0.80
Table 1: Cohen’s Kappa agreement between each pair of annotators, for each assessment question. Annotators were asked to edit the model response, if needed, in case they replied “yes” to the first two questions.
The high level of agreement observed confirmed the annotation task’s clarity and criteria consistency, allowing us to proceed with a single annotator annotating each dialog in the main phase.
@@ 4.2.1 Query assessment
Annotators verify query validity by answering the question: "Is the question understandable and makes sense?". If a query is marked as invalid, it is removed from the dialog. We also remove all subsequent dialog turns to avoid cascading errors. Valid queries are then annotated with metadata, including their type (‘general’, ‘specific’, or ‘yes/no’) and whether they depend on prior turns (‘context-dependent’ or not). Context-Dependent query templates specifically target context-dependent queries, so we expect these turns to be context dependent, but annotators still check for context-dependency to confirm that the query does depend on previous turns.
@@ 4.2.2 Response annotation: attribution, validation, and editing
Due to the current limitations of LLM-based attribution detection (which we also observed in §7.3.2 and is reported in the literature e.g. (Gao et al., 2023)), we opted for fully manual attribution annotation. For each generated response, annotators identify supporting text spans in the transcript while also verifying that the response is factually correct and editing the response as necessary to ensure its accuracy (more details in Appendix B). If a query cannot be answered from the transcript, it is marked as ‘unanswerable’ and annotators ensure the response accurately conveys the lack of information.
@@ 5 Datasets
This section describes the MISeD dataset (§5.1), constructed using our semi-authomatic method (§4). We also present a smaller independent dataset created using the fully manual Wizard-of-Oz methodology (§5.2), created to assess whether MISeD data is comparable to human-generated data. Finally, we present a manual assessment of MISeD response quality (§5.3).
@@ 5.2 Wizard-of-Oz Dataset
The collected WOZ data comprises 70 dialogs based on meetings from the test split, with a total of 700 query-response pairs. This data is used in subsequent evaluations (Section 7) for a comparison of model performance. Table 15 in the Appendix presents a comparison between the WOZ data and the MISeD data, with respect to response vocabulary and length.
@@
5.3 MISeD Response Quality Assessment
MISeD dialogs were generated by an LLM and then validated and corrected by humans (§4). To assess the quality of the agent responses created through this semi-automatic process, we aimed to compare them with responses created fully-manually. To that end, we collected human responses to a randomly selected sample of 100 MISeD queries. Annotators were provided with the meeting transcript and the dialog history up to the selected query and were tasked with answering the query.
Next, we provided a new group of annotators with the manually-generated responses, alongside the original MISeD response for each query (in randomized order between the two), and the corresponding query, dialog history, and meeting transcript. Annotators assessed which response was overall superior, considering correctness, grounding, and clarity, using a scale of ‘Equally Good’, ‘Slightly Better’, or ‘Substantially Better’. Table 3 presents the results.
Overall, MISeD responses were ranked as better in 55 pairs, compared to 43 pairs for fully-manual responses. A statistical sign test yields $P(x\geq 55)=0.13$ ($H_{0}:p=0.5$; $n=98$). These results suggest that the quality of MISeD responses is at least as good as fully-manual responses, and possibly somewhat better, though a larger sample would be needed to establish statistical significance.
Response evaluation categories Count Sum
MISeD ‘substantially better’ 29 55
MISeD ‘slightly better’ 26
‘equally good’ 2
Human-only ‘slightly better’ 21 43
Human-only ‘substantially better’ 22
@@ 6.1 Response Quality Evaluation
Agent responses are evaluated against the gold responses in the test dataset. Modern LLMs produce outputs that very often exhibit high readability and consistency, therefore we chose to focus our evaluation on the content.
@@ 6.1.1 Human evaluation
We conduct human evaluation on a random subset of 100 queries. The annotator is presented with the current user query, the gold response and the model responses, as well as the meeting transcript and the dialog history. To quantify the content overlap between the model response and the gold response, annotators provide scores for recall (how much of the gold response is covered by the model response), and precision (how much of the model response is covered by the gold response). Here covered refers to content overlap between the gold and model responses, in a semantic sense. Both scores are on a Likert scale of 1 to 4, corresponding to ‘Not Covered’, ‘Slightly Covered’, ‘Mostly Covered’ and ‘Fully Covered’.
@@ 6.2.2 Automatic evaluation
For automatic attribution evaluation we use the NLI model TrueTeacher (Gekhman et al., 2023) to approximate human entailment judgement, automatically computing recall (the proportion of response sentences fully entailed by the attributions) and precision (the proportion of attributions which contribute to the entailment of the response). Appendix C details the method implementation and discusses its limitations.
@@ 7 Baseline Models and Results
Response - Human scores Response - Automatic scores
@@
Attribution - Human scores Attribution - Automatic scores
@@ 7.1 Models
MISeD (n=2922) 1.91 1.97 29.73 10.90 23.64 0.40 0.28 0.25 0.26
MISeD (n=500) 1.93 1.99 30.52 11.10 24.10 0.40 0.26 0.16 0.20
WOZ (n=500) 1.91 1.94 31.88 12.64 26.17 0.41 0.02 0.36 0.04
Table 6: Evaluation using different training sets: the semi-automatic MISeD data and the WOZ manual data. Results on 200 WOZ test examples, using Gemini Pro finetuned.
@@ 7.3.1 Response quality
(3) The scores on WOZ test set exhibit the same trends, but are lower for all models. The lower performance for the non-finetuned models might suggest that fully-manual WOZ data is more challenging, being created freely by human annotators, compared to the templated methodology of MISeD. As expected, the finetuned models performed better on the MISeD test set, as it was created by the same protocol as the training data.
@@
Model Rouge-1 Rouge-2 Rouge-L BLEURT
Gemini Pro 27.82 10.08 22.03 0.38
Gemini Pro Finetuned on MISeD 30.31 11.39 24.26 0.40
Gemini Pro Finetuned on QMSum 27.72 8.89 21.60 0.38
Table 8: Average scores for response evaluation on the WOZ test sets.
@@ 7.3.2 Attribution quality
Observing the more reliable human scores, we see that the zero-shot models suffer from low recall, due to their tendency to provide attribution less frequently. Finetuning with MISeD data improves performance, substantially increasing recall with a small drop in precision, notably increasing F1 scores for both test sets. The automatic scores show similar though weaker trends, but we regard them as less reliable (see Appendix C).
@@ 7.4 Comparison of Training Sets
In this section we compare the performance of models trained on our semi-automatic MISeD data to those trained on the manual WOZ data. We split our WOZ dataset (700 dialog turns) into training and test sets (500:200, respectively). We compare training with (i) 500 WOZ examples (ii) 500 semi-automatic MISeD examples (iii) the full MISeD training set (2922 examples). For response generation, Table 6 suggests that training on 500 semi-automatic MISeD examples yields comparable results to training on the same amount of fully manual examples. For attribution, the model benefits from a larger number of training examples, suggesting that 500 examples are insufficient for this subtask. Notably, the model trained on WOZ only predicted attributions for 6 out of 200 test examples; thus we didn’t perform a manual attribution evaluation (as in Table 5). We also tried adding the WOZ data to the MiSeD data in finetuning, but, with respect to automatic evaluation, this did not improve notably the response quality while attribution quality was notably deteriorated, hence we did not pursue this data combination further. To summarize, training on our semi-automatic MISeD data yields response generation quality that is on par with training on the fully manual WOZ data, while also improving attribution, and saving time and effort.
@@ Appendix B Annotation details
Appendix C Automatic Evaluation of Attribution Implementation
@@
WOZ MISeD Test Set
# meetings 35 34
# dialogs 70 66
# query-response pairs 700 628
overall vocabulary size 2012 2492
avg. response length 28.52 41.71
avg. response vocabulary size 23.18 30.53
avg. % of response vocabulary overlapping with the transcript 84.86% 78.43%
Table 15: Statistics of the WOZ test set compared to the MISeD test set. MISeD responses tend to be longer, and to use a broader vocabulary, overlapping less with the transcript vocabulary. Vocabulary size is calculated as the number of unique lemmas.